OpenStudy (anonymous):

You have a committee of 13 members, made up of 8 men and 3 women. In how many ways can 3 tasks be assigned so that both men and women are given assignments?

OpenStudy (anonymous):

The number of ways in which 3 tasks can be distributed such that only men or only women get the job is. Only men 8C3 Only women : 3C3=1 Total ways: 13C3 Hence the answer would be 13C3-1-8C3

OpenStudy (anonymous):

ill work it out

OpenStudy (anonymous):

that wasnt right

OpenStudy (anonymous):

Hmm okay maybe it was wrong i am not sure....

OpenStudy (anonymous):

would it be a permutation

OpenStudy (anonymous):

No i think it'll be a combination only.

OpenStudy (anonymous):

One task is given to one person only right?

OpenStudy (anonymous):

it just says how many ways can the 3 tasks be assigned so that both women and men are given assignments

OpenStudy (dumbcow):

is that a typo? shouldn't it be 8 men and 5 women

OpenStudy (anonymous):

total 13 members and 3 total tasks

OpenStudy (anonymous):

Is 220 the answer?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

shankvee had the right idea :) but your numbers don't add up ....8+3 = 11 not 13

OpenStudy (anonymous):

leave u cant read 8+5

OpenStudy (dumbcow):

?? 13C3 - 8C3 - 5C3 = 286 - 56 - 10 = 220

OpenStudy (anonymous):

sorry thought you were messing with me i am so sorry but 220 is not right

OpenStudy (dumbcow):

maybe it is permutations then.... 13P3 - 8P3 - 5P3 = 1320

OpenStudy (anonymous):

you are a genius

OpenStudy (anonymous):

It was permutations

OpenStudy (anonymous):

I am so sorry for my comment earlier I had some punk on here earlier and thought you were messing with me

OpenStudy (anonymous):

I am very very sorry

OpenStudy (dumbcow):

no prob, just noticed a typo in your question post

OpenStudy (anonymous):

I am so sorry

OpenStudy (anonymous):

I have tried the problem 40 times

OpenStudy (dumbcow):

do you know the formulas for combinations and permutations? most calculators will do them for you as well

OpenStudy (anonymous):

somewhat...i find them confusing and hard to remember also my finite course does not allow calculators during exams so i am trying to do them manually because all of our test questions come from these particular set of problems

OpenStudy (anonymous):

I have a few more permuations and maybe a combo i am going to post that are two part questions....3-5 of these have stumped me, most i can work through

Directrix (directrix):

Should this 13 be 11? Or, are the other two committee members neither female nor male? Just wondering. 13 members, made up of 8 men and 3 women. 8+3 =11. I don't think this is a permutation problem. I could be wrong but what about the 13 committee - those two extra members?